OpenStudy (anonymous):

Suppose you have 50 feet of fencing to enclose a rectangular dog pen. The function A=25x-x^2, where x=width, gives you the area of the dog pen in square feet. What width gives you the maximum area? What is the maximum are?

OpenStudy (anonymous):

maximum is the second coordinate of the vertex the first coordinate of the vertex is always \(-\frac{b}{2a}\) which in your case is \(-\frac{25}{2\times -1}=12.5\)

OpenStudy (anonymous):

so your final answer is 12.5?

OpenStudy (anonymous):

in other words, in plain english, build a square

OpenStudy (anonymous):

your question has two parts: a) find the width, and b) find the area

OpenStudy (anonymous):

huh? I got A=25-x^2 using -b/2a and got -50/2-1 and got width equaling 25

OpenStudy (anonymous):

yes that is the question

OpenStudy (anonymous):

the answer to the first part is the width and the length should both be \(12.5\)

OpenStudy (anonymous):

\(b=25\) not \(50\)

OpenStudy (anonymous):

so the maxium are would be.....

OpenStudy (anonymous):

12.5x12.5?

OpenStudy (anonymous):

whatever \(12.5\times 12.5\) is or whatever \(25\times 12.5-12.5^2\) is either way, it is the same number

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

ok i have one more problem for you

OpenStudy (anonymous):

it is fairly easy

OpenStudy (anonymous):

ru up for it

terenzreignz (terenzreignz):

What is it?